OpenStudy (anonymous):

∫csc5xcot5x dx please integrate :) tnx

OpenStudy (anonymous):

I would try using some trig identities on that.

OpenStudy (anonymous):

yes that our topic about trigonometric identities . ill wait for your reply thank you

OpenStudy (anonymous):

Remember \[\csc x = 1/\sin x\] and \[\cot x = \cos x/ \sin x\]

OpenStudy (anonymous):

d/dx[cscx] = -cscx*cotx

OpenStudy (anonymous):

@erish23 I'm not going to do your problem for you. That's the part of learning, I give you hints, you figure it out. :)

OpenStudy (anonymous):

ah ok . thank you :) i know how to solve it thank you for your help :)